# Variational Mode-Driven Graph Convolutional Network for Spatiotemporal Traffic Forecasting

### v1

@@ II-C Problem Formulation
The dynamic evolving graph can be defined as time-dependent $\mathcal{G}^{(t)}=(V^{(t)},A^{(t)},X^{(t)})$. We assume that the vertices are constant but the features are time-dependent and are also influenced by unforeseen conditions such as road accidents, road closures, and public events, which alter the course of traffic flow. In this work, we aim to predict the future horizon traffic flow [$\boldsymbol{X}^{(T_{w}+1)},\boldsymbol{X}^{(T_{w}+2)},\dots,\boldsymbol{X}^{(T_{w}+N_{H})}$] for the next prediction horizon length $N_{H}$ given $\mathcal{G}^{(t)}$ for $t\in[1,T]$, i.e.,
$\begin{split}[\mathcal{G}^{(1)},\dots,\mathcal{G}^{(T_{w})}]\stackrel{{\scriptstyle F}}{{\rightarrow}}[\boldsymbol{X}^{(T_{w}+1)},\dots,\boldsymbol{X}^{(T_{w}+N_{H})}],\end{split}$
@@
where $T$ is the length of the signal and $F$ is a deep-learning-based function that takes current and historical TTNG as input to yield the future TTNG.
III Methodology
III-A Mathematical Formulation
@@
The mathematical formulation of the proposed framework is divided into two main steps. The first step focuses on the pre-processing of the time signal using variational mode decomposition (VMD) [31]. The second step involves the deep-learning model for spatio-temporal forecasting [11].
III-A1 Variational Mode Decomposition (VMD)
@@
The objective of VMD is to decouple the original signal into sub-signals $u_{k}$, known as modes. The analytic signal $f_{A}(t)$ is a complex-valued function that contains only non-negative frequency components. An oscillatory signal is a valid intrinsic mode function (IMF) if it fulfills the two main conditions; the number of local extrema (minima and maxima) and zero-crossings that differ at most by one; the average value of the mode is zero [27]. The analytic signal is a complex signal where the real part comprises the original signal $f(t)$ and the imaginary part is the Hilbert transform $\mathcal{H}f(t)$ of $f(t)$, that is
$f_{A}(t)=f(t)+j\mathcal{H}f(t).$ (1)
@@
The analytic signal is used for envelope detection in the oscillatory signal. We solve the following optimization problem to decompose the analytic signal $f_{A}(t)$ into its variational modes [31]:
@@
$\min_{u_{k},\omega_{k}}\quad&\sum_{k}\lVert\partial_{t}[f_{A}(t)*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2},\\ &\text{such that}\quad\sum_{k}u_{k}=f,\end{aligned}$ (2)
@@
Here, $u_{k}=\{u_{1},u_{2},\dots,u_{K}\}$ and $\omega_{k}=\{\omega_{1},\omega_{2},\dots,\omega_{K}\}$ describes sets of all $K$ modes and their center frequencies, respectively. If the Dirac delta function is the original function $f=\delta(t)$, the impulse response of the Hilbert transform will be $\frac{j}{\pi t}$. The objective function in (2) can be expressed as
$\min_{u_{k},\omega_{k}}&\left\{\sum_{k}\lVert\partial_{t}[(\delta(t)+\frac{j}{\pi t})*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2}\right\},\\ &\text{such that}\quad\sum_{k}u_{k}=f.\end{aligned}$ (3)
@@
$\begin{split}\mathcal{L}(u_{k},\omega_{k},\lambda)=\alpha\sum_{k}\lVert\partial_{t}[(\delta(t)+\frac{j}{\pi t})*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2}+\\ \lVert f(t)-\sum_{k}u_{k}(t)\rVert_{2}^{2}+\langle\lambda(t),f(t)-\sum_{k}u_{k}(t)\rangle,\end{split}$ (4)
$u_{k}^{(n+1)}=\rm{arg}\min_{u_{k}}\mathcal{L}(\{u^{n+1}_{i<k},u^{n}_{i\geq k}\},\omega_{i}^{n},\lambda^{n}),$ (5)
$\omega_{k}^{(n+1)}=\rm{arg}\min_{\omega_{k}}\mathcal{L}(u^{n+1}_{i},\{\omega^{n+1}_{i<k},\omega^{n}_{i\geq k}\},\lambda^{n}),$ (6)
where $\alpha$ is the penalty variable in the variational mode term. The second term is a quadratic penalty for the reconstruction fidelity term. This term ensures the convergence properties at finite weight. The Lagrangian multipliers $\lambda(t)$ strictly enforce the constraints. The solution of the mode minimization problem is given by
$\hat{u}_{k}^{(n+1)}(\omega)=\frac{\hat{f}(\omega)-\sum\limits_{i\neq k}\hat{u_{i}}(\omega)+\frac{\hat{\lambda}(\omega)}{2}}{1+2\alpha(\omega-\omega_{k})^{2}},$ (7)
where $\alpha$ is a regularization term to control the bandwidth of the modes. It is a trade-off between the accuracy and the smoothness of the bandwidth of the decomposed modes. The formulation of the minimization problem for center frequency (6) is given by
@@
The above optimization problems consist of minimization of two variables $\hat{u}_{k}(w)$ and $\omega_{k}$. To solve this, the alternating direction method of multipliers (ADMM) is an optimization technique that decomposes this dual optimization in two sub-problems and solves each sub-problem individually as described in Algorithm 1.
@@
     $\hat{u}_{k}^{(n+1)}=\frac{\hat{f}(\omega)-\sum\limits_{i<k}\hat{u_{i}}^{n+1}(\omega)-\sum\limits_{i>k}\hat{u_{i}}^{n}(\omega)+\frac{\hat{\lambda}(\omega)}{2}}{1+2\alpha(\omega-\omega^{n}_{k})^{2}}$
@@
     $\omega_{k}^{n+1}=\frac{\sum\limits_{\omega=T}^{2T}\omega|\hat{u}_{k}^{n+1}(\omega)|^{2}d\omega}{\sum\limits_{\omega=T}^{2T}|\hat{u}_{k}^{n+1}(\omega)|^{2}d\omega}$
@@
The pre-processing of VMD is also a crucial step in this optimization process. Since the input of the signal is in the time domain consisting of finite $T_{w}\in[1,T]$ and discrete samples, we have to convert it into a frequency domain signal using the discrete Fourier transform (DFT). The implicit assumption is that the signal is considered to be periodic in nature representing a one-time period. The discontinuity occurs at the endpoints of the time signal. To resolve this discontinuity, the time signal is mirrored on both ends around the central axis before applying DFT.
The post-processing of VMD is the inverse operation where the frequency domain signal is converted back to the time domain signal using inverse discrete Fourier transform (IDFT). The truncation operation is applied for the reversal operation of the mirror to convert the signal to the original length.
$\phi(t)=f(t)-\sum_{i=0}^{K}u_{i}(t),$
where redemption function $\phi(t)$ is the difference between the original signal and the summation of the modes at time $t$.
III-A2 Attention Based Spatial-Temporal Graph Convolutional Networks (ASTGCN)
The graph feature within a time window $T_{w}$ is defined as the $\mathbf{\mathcal{X}}=[\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{T}]\in\mathbb{R}^{N\times d\times T_{w}}$, where $\mathbf{X}_{1}\in\mathbb{R}^{d\times T_{w}}$ is the feature matrix for the graph at $t=1$. The number of feature channels is denoted by $d$. The spatial attention on the graph feature matrix is given by:
$\mathbf{S}=\mathbf{V}_{s}\sigma((\mathbf{\mathcal{X}}\mathbf{W}_{1}\mathbf{W}_{2}(\mathbf{W}_{3}\mathbf{\mathcal{X}})^{\textit{T}})+\mathbf{b}_{s}),$ (9)
$\mathbf{S}^{{}^{\prime}}_{i,j}=\frac{\exp(\mathbf{S}_{i,j})}{\sum\limits_{j=1}^{N}\exp(\mathbf{S}_{i,j})},$ (10)
where $\mathbf{V}_{s},\>\mathbf{b}_{s}\in\mathbb{R}^{N\times N},\>\mathbf{W}_{1}\in\mathbb{R}^{T_{w}},\>\mathbf{W}_{2}\in\mathbb{R}^{d\times T_{w}},\>\mathbf{W}_{3}\in\mathbb{R}^{d}$ are learnable weights characterizing the spatial attention. $\sigma$ is the sigmoid activation function. $\mathbf{S}^{{}^{\prime}}_{i,j}$ is the normalized spatial correlation matrix between node $i$ and node $j$, which is dynamically computed. The temporal attention is described by the following expression:
$\mathbf{E}=\mathbf{V}_{e}\sigma(((\mathbf{\mathcal{X}}^{\textit{T}}\mathbf{U}_{1})\mathbf{U}_{2}(\mathbf{U}_{3}\mathbf{\mathcal{X}}))+\mathbf{b}_{e}),$ (11)
$\mathbf{E}^{{}^{\prime}}_{i,j}=\frac{\exp(\mathbf{E}_{i,j})}{\sum\limits_{j=1}^{T_{w}}\exp(\mathbf{E}_{i,j})},$ (12)
where $\mathbf{V}_{e},\>\mathbf{b}_{e}\in\mathbb{R}^{T_{w}\times T_{w}},\>\mathbf{U}_{1}\in\mathbb{R}^{N},\>\mathbf{U}_{2}\in\mathbb{R}^{N\times d},\>\mathbf{U}_{3}\in\mathbb{R}^{d}$ are trainable parameters to determine the temporal attention. $\mathbf{E}^{{}^{\prime}}_{i,j}$ is the normalized temporal correlation matrix between node $i$ and node $j$. In spectral graph analysis, graph connectivity is examined using the properties associated with the adjacency matrix or Laplacian matrix. The adjacency matrix for the weighted graph is defined as:
@@ III-A2 Attention Based Spatial-Temporal Graph Convolutional Networks (ASTGCN)
where $l_{i,j}$ is the road network distance between node $i$ and $j$, $\sigma$ denotes the standard deviation of distances, and $r$ is the threshold. The Laplacian matrix is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, $\mathbf{D}\in\mathbb{R}^{N\times N}$ is the diagonal matrix consisting of the degree of each node $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{(i,j)}$, and the normalized Laplacian matrix is expressed as $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. The eigenvalue decomposition of the Laplacian matrix is $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\textit{T}}$, where eigenvalues $\mathbf{\Lambda}={\rm diag}([\lambda_{1},\lambda_{2},\dots,\lambda_{N-1}])$ is a diagonal matrix. In graph Fourier transform, eigenvectors $\mathbf{U}$ form a Fourier basis consisting of complex sinusoids, and $\mathbf{U}$ is an orthogonal matrix. The kernel $g_{\theta}$ with the parameters $\theta$ is applied on the Laplacian matrix given by
$g_{\theta}(\mathbf{L})=g_{\theta}(\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\textit{T}})=\mathbf{U}g_{\theta}(\mathbf{\Lambda})\mathbf{U}^{\textit{T}}x.$
@@
The eigenvalue decomposition in spectral graphs is a computationally expensive task. Therefore, the filtering operation in the graph spectral domain can be formulated using the Chebyshev graph convolution network [38]:
$g_{\theta}(\mathbf{L})=\sum_{m=0}^{M-1}\theta_{m}(T_{m}(\hat{\mathbf{L}})\odot\mathbf{S}^{{}^{\prime}})x,$ (13)
@@
where $\theta$ is a Chebyshev coefficients and $T_{m}(\hat{\mathbf{L}})$ is the Chebyshev polynomial of order $m$. $\hat{\mathbf{L}}=2\mathbf{L}/\lambda_{\rm{max}}-I_{n},\lambda_{\rm{max}}$ is the maximum eigenvalue of the Laplacian matrix.
Figure 3: Block diagram of the variational mode graph convolutional network (VMGCN). The data is decomposed into components using VMD which are then concatenated as the features along the time of the day and day of the week. The original signal can also be concatenated in architecture variants and pass these modified features graph into a neural network. A spatio-temporal (ST) block applies SAtt: Spatial attention, TAtt: temporal attention, graph convolution, and time convolution. The prediction outcome $\hat{\mathcal{X}}\in\mathbb{R}^{N\times T_{w}}$ of VMGCN contains the future traffic flow from the timestamp of ($T_{w}+1$) to ($T_{w}+N_{H}$).
III-B Architecture Detail

### v3

@@ 2.3. Problem Formulation
The dynamic evolving graph can be defined as time-dependent $\mathcal{G}^{(t)}=(V^{(t)},\,A^{(t)},\,Y^{(t)})$. We assume that the vertices are constant, but the features are time-dependent and are also influenced by unforeseen conditions such as road accidents, road closures, weather, and public events, which alter the course of traffic flow. In this work, we aim to predict the future horizon traffic flow [$Y^{(t+1)},\,Y^{(t+2)},\dots,\,\boldsymbol{Y}^{(t+N_{H})}$] for the next prediction horizon length $N_{H}$ given $\mathcal{G}^{(t)}$ for $t\in[1,\,T-N_{H}]$, i.e.,
$\begin{split}[Y^{(t-T_{w}+1)},\dots,Y^{(t)}]\stackrel{{\scriptstyle F}}{{\rightarrow}}[Y^{(t+1)},\dots,Y^{(t+N_{H})}],\end{split}$
@@
where $T$ is the length of the signal, $N_{H}$ is the length future horizon, and $F$ is a deep-learning-based function that takes current and historical TTNG as input to yield the future TTNG.
3. Methodology
3.1. Mathematical Formulation
@@
The proposed prediction framework is comprised of pre-processing of the time signal using variational mode decomposition (VMD) (Dragomiretskiy and Zosso, 2013) followed by the spatiotemporal forecasting using the deep-learning model (Guo et al., 2019).
3.1.1. Variational Mode Decomposition (VMD)
@@
The objective of VMD is to decouple the original signal into sub-signals $u_{k}$, known as modes. The analytic signal $f_{A}(t)$ is a complex-valued function that contains only non-negative frequency components. An oscillatory signal is a valid intrinsic mode function (IMF) if it fulfills the two main conditions; (1) the number of extrema (minima and maxima) and zero-crossings must either equal or differ at most by one; (2) at any point, the mean value of the envelope defined by the local maxima and the envelope defined by the local minima is zero (as defined in (Huang et al., 1998)). The analytic signal is a complex signal where the real part comprises the original signal $f(t)$ and the imaginary part is the Hilbert transform $\mathcal{H}f(t)$ of $f(t)$, that is
(1) $f_{A}(t)=f(t)+j\mathcal{H}f(t).$
@@
We solve the following optimization problem to decompose the analytic signal $f_{A}(t)$ into its variational modes (Dragomiretskiy and Zosso, 2013):
@@
(2) $\min_{u_{k},\omega_{k}}\quad&\sum_{k}\lVert\partial_{t}[f_{A}(t)*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2},\\ &\text{such that}\quad\sum_{k}u_{k}=f,\end{aligned}$
@@
Here, $u_{k}=\{u_{1},u_{2},\dots,u_{K}\}$ and $\omega_{k}=\{\omega_{1},\omega_{2},\dots,\omega_{K}\}$ are sets of all $K$ modes and their center frequencies, respectively. If the Dirac delta function is the original function $f=\delta(t)$, the impulse response of the Hilbert transform will be $\frac{j}{\pi t}$. The objective function in (2) can be expressed as
(3) $\min_{u_{k},\omega_{k}}&\left\{\sum_{k}\lVert\partial_{t}[(\delta(t)+\frac{j}{\pi t})*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2}\right\},\\ &\text{such that}\quad\sum_{k}u_{k}=f.\end{aligned}$
@@
(4) $\begin{split}\mathcal{L}(u_{k},\omega_{k},\lambda)=\alpha\sum_{k}\lVert\partial_{t}[(\delta(t)+\frac{j}{\pi t})*u_{k}(t)]e^{-j\omega_{k}t}\rVert_{2}^{2}+\\ \lVert f(t)-\sum_{k}u_{k}(t)\rVert_{2}^{2}+\langle\lambda(t),f(t)-\sum_{k}u_{k}(t)\rangle,\end{split}$
(5) $u_{k}^{(n+1)}=\rm{arg}\min_{u_{k}}\mathcal{L}(\{u^{n+1}_{i<k},u^{n}_{i\geq k}\},\omega_{i}^{n},\lambda^{n}),$
(6) $\omega_{k}^{(n+1)}=\rm{arg}\min_{\omega_{k}}\mathcal{L}(u^{n+1}_{i},\{\omega^{n+1}_{i<k},\omega^{n}_{i\geq k}\},\lambda^{n}),$
where $\alpha$ is the bandwidth constraint in the variational mode term to control the bandwidth of the modes. The second term is a quadratic penalty for the reconstruction fidelity term and ensures the convergence properties at finite weight. The Lagrangian multipliers $\lambda(t)$ strictly enforce the constraints. The solution to the mode minimization problem is given by
(7) $\hat{u}_{k}^{(n+1)}(\omega)=\frac{\hat{f}(\omega)-\sum\limits_{i\neq k}\hat{u_{i}}(\omega)+\frac{\hat{\lambda}^{n}(\omega)}{2}}{1+2\alpha(\omega-\omega_{k}^{n})^{2}},$
where $\hat{f}(\omega)$ is the Fourier Transform of $f$ and $\hat{u}_{k}(\omega)$ denotes the modes in frequency domain. By adjusting the bandwidth constraint, we trade off between the accuracy and smoothness of the bandwidth of the decomposed modes. The formulation of the minimization problem for center frequency (6) is given by
@@
To solve the above optimization problems that require the minimization of two variables $\hat{u}_{k}(\omega)$ and $\omega_{k}$, the alternating direction method of multipliers (ADMM) decomposes this dual optimization in two sub-problems and solves each sub-problem individually as described in Algorithm 1. Since the input of the signal is in the time domain, consisting of finite and discrete samples, the pre-processing step involves the use of discrete Fourier transform (DFT) to obtain a frequency domain signal. The implicit assumption is that the signal, representing a one-time period, is periodic in time. This assumption results in a discontinuity at the endpoints which can be resolved by mirroring the time signal on both ends around the central axis before applying DFT. The post-processing involves the inverse operation, where the frequency domain signal is converted back to the time domain signal using the inverse DFT. The truncation of the signal is carried out to reverse the mirroring operation to convert the signal to the original length.
$\phi(t)=f(t)-\sum_{k=1}^{K}u_{k}(t),$
where the redemption (residual) function $\phi(t)$ is the difference between the original signal and the summation of the modes at time $t$.
@@
     $\hat{u}_{k}^{(n+1)}=\frac{\hat{f}(\omega)-\sum\limits_{i<k}\hat{u_{i}}^{n+1}(\omega)-\sum\limits_{i>k}\hat{u_{i}}^{n}(\omega)+\frac{\hat{\lambda}^{n}(\omega)}{2}}{1+2\alpha(\omega-\omega^{n}_{k})^{2}}$
@@
     $\omega_{k}^{n+1}=\frac{\sum\limits_{\omega=T}^{2T}\omega|\hat{u}_{k}^{n+1}(\omega)|^{2}}{\sum\limits_{\omega=T}^{2T}|\hat{u}_{k}^{n+1}(\omega)|^{2}}$
@@
Figure 3. Block diagram of the variational mode graph convolutional network (VMGCN). The data is decomposed into components using VMD which are then concatenated as the features along the time of the day and day of the week. The original signal can also be concatenated in architecture variants and pass these modified features graph into our attention-augmented GCN. A spatiotemporal (ST) block applies SAtt: Spatial attention, TAtt: temporal attention, graph convolution, and time convolution. The prediction outcome $\mathcal{X}\in\mathbb{R}^{N\times T_{w}}$ of VMGCN contains the future traffic flow from the timestamp of ($T_{w}+1$) to ($T_{w}+N_{H}$).
3.1.2. Attention-Based Spatial-Temporal Graph Convolutional Networks (ASTGCN)
The input graph feature within a time window $T_{w}$ is defined as the $\mathbf{\mathcal{X}}=[\mathbf{X}_{t-T_{w}+1},\mathbf{X}_{t-T_{w}+2},\dots,\mathbf{X}_{t}]\in\mathbb{R}^{N\times d\times T_{w}}$, where $\mathbf{X}_{1}\in\mathbb{R}^{d\times T_{w}}$ is the feature matrix for the graph at $t=1$. The number of feature channels is denoted by $d$. The spatial attention on the graph feature matrix is given by
(9) $\mathbf{S}=\mathbf{V}_{s}\sigma((\mathbf{\mathcal{X}}\mathbf{W}_{1}\mathbf{W}_{2}(\mathbf{W}_{3}\mathbf{\mathcal{X}})^{\textit{T}})+\mathbf{b}_{s}),$
(10) $\mathbf{S}^{{}^{\prime}}_{i,j}=\frac{\exp(\mathbf{S}_{i,j})}{\sum\limits_{j=1}^{N}\exp(\mathbf{S}_{i,j})},$
where $\mathbf{V}_{s},\>\mathbf{b}_{s}\in\mathbb{R}^{N\times N},\>\mathbf{W}_{1}\in\mathbb{R}^{T_{w}},\>\mathbf{W}_{2}\in\mathbb{R}^{d\times T_{w}},\>\mathbf{W}_{3}\in\mathbb{R}^{d}$ are learnable weights characterizing the spatial attention. $\sigma$ is the sigmoid activation function. $\mathbf{S}^{{}^{\prime}}_{i,j}$ is the normalized spatial correlation matrix between node $i$ and node $j$, which is dynamically computed. The temporal attention is described as
(11) $\mathbf{E}=\mathbf{V}_{e}\sigma(((\mathbf{\mathcal{X}}^{\textit{T}}\mathbf{U}_{1})\mathbf{U}_{2}(\mathbf{U}_{3}\mathbf{\mathcal{X}}))+\mathbf{b}_{e}),$
(12) $\mathbf{E}^{{}^{\prime}}_{i,j}=\frac{\exp(\mathbf{E}_{i,j})}{\sum\limits_{j=1}^{T_{w}}\exp(\mathbf{E}_{i,j})},$
where $\mathbf{V}_{e},\>\mathbf{b}_{e}\in\mathbb{R}^{T_{w}\times T_{w}},\>\mathbf{U}_{1}\in\mathbb{R}^{N},\>\mathbf{U}_{2}\in\mathbb{R}^{N\times d},\>\mathbf{U}_{3}\in\mathbb{R}^{d}$ are trainable parameters to determine the temporal attention. $\mathbf{E}^{{}^{\prime}}_{i,j}$ is the normalized temporal correlation matrix between node $i$ and node $j$. In spectral graph analysis, graph connectivity is examined using the properties associated with the adjacency matrix or Laplacian matrix. The binary adjacency matrix for the weighted graph is defined by using a Gaussian kernel thresholded function (Shuman et al., 2013)
@@ 3.1.2. Attention-Based Spatial-Temporal Graph Convolutional Networks (ASTGCN)
where $l_{i,j}$ is the road network distance between node $i$ and $j$, $\sigma$ denotes the standard deviation of distances, and $r$ is the threshold. The Laplacian matrix is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, $\mathbf{D}\in\mathbb{R}^{N\times N}$ is the diagonal matrix consisting of the degree of each node $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{(i,j)}$, and the normalized Laplacian matrix is expressed as $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. The eigenvalue decomposition of the Laplacian matrix is $\mathbf{L}=\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{\textit{T}}$, where eigenvalues $\mathbf{\Lambda}={\rm diag}([\lambda_{1},\lambda_{2},\dots,\lambda_{N-1}])$ is a diagonal matrix. In graph Fourier transform, eigenvectors $\mathbf{Q}$ form a Fourier basis consisting of complex sinusoids, and $\mathbf{Q}$ is an orthogonal matrix. The kernel $g_{\theta}$ with the parameters $\theta$ is applied on the Laplacian matrix given by
$g_{\theta}(\mathbf{L})=g_{\theta}(\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{\textit{T}})=\mathbf{Q}g_{\theta}(\mathbf{\Lambda})\mathbf{Q}^{\textit{T}}x.$
@@
Since the eigenvalue decomposition in spectral graphs is a computationally expensive task, the filtering operation in the graph spectral domain can be formulated using the Chebyshev graph convolution network as (Defferrard et al., 2016)
(13) $g_{\theta}(\mathbf{L})=\sum_{m=0}^{M-1}\theta_{m}(T_{m}(\hat{\mathbf{L}})\odot\mathbf{S}^{{}^{\prime}})\mathcal{X},$
@@
where $\theta$ is a Chebyshev coefficients, $T_{m}(\hat{\mathbf{L}})$ is the Chebyshev polynomial of order $m$, and $\hat{\mathbf{L}}=2\mathbf{L}/\lambda_{\rm{max}}-I_{n}$. Here, $\lambda_{\rm{max}}$ is the maximum eigenvalue of the Laplacian matrix $\mathbf{L}$.
3.2. Model Architecture